# Climate-economy projections under shared socioeconomic pathways and net-zero scenarios

### v1

@@ 2.2 DICE calibration
Population $L_{t}$, total factor productivity $A_{t}$, land emission $E_{t}^{\mathrm{Land}}$, and decorbanisation $\sigma_{t}$ are exogenous parametric functions specified in DICE. In the original DICE, these were estimated using various data sources, resulting in a single scenario of the future for the economy and climate. In this study we estimate these exogenous functions from population, GDP and emission data of the baseline SSPs and process-based IAMs, i.e. we get many plausible scenarios that allow to assess the possible trajectories of climate and economic variables in the future.
The SSPs provide quantitative projections and qualitative descriptions. Quantitative projections include population and economic growth. A detailed description of SSPs can be found in Riahi et al. (2017). In short, they are referred as:
  • SSP1 – “Sustainability - taking the Green Road” (low challenges to mitigation and adaptation), representing the sustainable development;
  • SSP2 – “Middle of the Road” (medium challenges to mitigation and adaptation), implies a development pathway consistent with typical historical patterns;
  • SSP3 – “Regional Rivalry – A Rocky Road” (high challenges to mitigation and adaptation), characterized by international fragmentation and regional rivalry;
  • SSP4 – “Inequality – A Road Divided (low challenges to mitigation, high challenges to adaptation)”, featuring low challenges to mitigation but high challenges to adaptation emphasizing extreme inequality;
  • SSP5 – “Fossil-fueled Development – Taking the Highway” (high challenges to mitigation, low challenges to adaptation), representing high challenges to mitigation and low challenges to adaptation.
@@
There are many process-based IAMs. These are detailed, multi-region, large-scale computer simulation models to assesses policy options for stabilizing the global climate. They couple representations of the economy, the energy system, the agricultural and land use system, and the climate system. The specific IAMs considered in our study will be described below.
In principle, each process-based IAM can be run for each SSP. However, for reference, for each SSP a single IAM interpretation was selected as the so-called representative marker scenario for recommended use by future analyses of climate change, its impacts, and response measures. These are the so-called marker SSPs selected as representative of the broader developments of each SSP, guided by the internal consistency and the ability of the different models to represent distinct characteristics of the storylines. Below is a summary list of marker SSPs:
@@
For a detailed discussion, see also Riahi et al. (2017). Figure 1 shows the data for GDP, population and emission under five marker SSPs (baselines). Here, we note that only under SSP3, world population keeps growing from now to 2100. Under other SSPs, the population peaks at around 2050-2070. GDP grows over time under all SSPs and is smallest under SSP3. Industry emission grows under SSP2 and SSP3. It increases until 2050-2060 for SSP1 and SSP4, and then decreases. Under SSP5, emission is the largest, growing fast until around 2080 where it flattens. Under SSP1, emission is the lowest.
Figure 1: Global population (millions), GDP (trillions 2005 USD per year), and industrial $\mathrm{CO}_{2}$ emission (Gigaton) scenarios by marker SSPs (Red: SSP1, Grey: SSP2, Blue: SSP3, Green: SSP4, Brown: SSP5).
The data are available from the SSPdata-IIASA database. In addition to data from marker SSPs, we also considered data for non-marker SSPs. Not all IAMs can run for all SSPs. For non-marker scenarios, we consider four IAMs: AIM/CGE, GCAM, IMAGE and WITCH-GLOBIOM Emmerling et al. (2016) that have results under all five SSPs.
We use the baseline SSP scenarios from the process-based IAMs to estimate the DICE exogenous functions as follows (the corresponding variables taken from the SSP-IAM scenario are indicated by $\widehat{}$ ).
  • The population $L_{t}$ is taken from the SSP-IAM scenario, i.e. $L_{t}=\widehat{L}_{t}$.
  • The decarbonisation function $\sigma_{t}$ is estimated as $\sigma_{t}=\widehat{E}^{\mathrm{Ind}}_{t}$/$\widehat{Y}_{t}$, where $\widehat{E}^{\mathrm{Ind}}_{t}$ is taken from the IAM baseline scenarios and $\widehat{Y}_{t}$ is the gross GDP in the baseline SSPs.
@@
  • The total productivity factor $A_{t}$ is calibrated so that the gross GDP $Y_{t}$ from DICE matches $\widehat{Y}_{t}$ the gross GDP in the baseline SSPs. This calibration involves iterative process described in Yang et al. (2018) that, for each iteration $m=1,2,...$ calculates the DICE trajectories of $Y^{(m)}_{t}$ and model state variables $X^{(m)}_{t}$, $t=0,1,...,N$ using
@@
    until trajectory of $Y^{(m)}_{t}$ approximates the SSP trajectory of $\widehat{Y}_{t}$ closely. Each iteration involves solving the DICE model (optimizing the welfare function with respect to the consumption) without climate damage and abatement costs, $K^{(0)}_{t}$ is taken from the original DICE solution, and convergence is achieved after few iterations.
Then for each baseline SSP and process-based IAM we solve DICE (miximising the welfare function wrt consumption $C_{t}$) resulting in determinstic trajectory for temperature, carbon concentration and social cost of carbon. We do not optimise welfare function in DICE for emission control $\widetilde{\mu}_{t}\in[0,1]$ but we consider scenarios when net-zero emission is achieved by a specific year as follows:
  • Net-zero in 2050: $\tilde{\mu}_{t}$ increases linearly from 0 in 2015 to 1 in the year 2050 and remains 1 afterward; and
  • Net-zero in 2100: $\tilde{\mu}_{t}$ increases linearly from 0 in 2015 to 1 in the year 2100 and remains 1 afterward.
@@ 3 Results
Figure 2 shows the DICE results for the trajectories of atmospheric temperature, atmospheric concentration and social cost of carbon under five SSP markers from 2015 to 2100. Each subplot in Figure 2 shows the trajectory under the original DICE (black dashed line), the trajectory under the scenario of achieving net-zero emissions in 2050 (red line) and the trajectory under the scenario of achieving net-zero emission in 2100 (blue line). Here, the original DICE trajectory is for benchmarking.
As expected we can see that the temperature under the scenario of net-zero in 2050 is lower than under the scenario of net-zero in 2100 across all SSPs. Under the 2050 net-zero scenario, the temperature in 2100 is about 2.5C for all SSPs while under the 2100 net-zero scenario, temperature in 2100 is just under 3C for SSP1 and materially larger than 3C for SSP2-SSP5. This is of course not surprising because SSP1 corresponds to “Green road” scenario. The original DICE trajectory of temperature is close 2100 net-zero scenario. This is somewhat expected because the optimal $\mu_{t}$ under original DICE reaches 1 in 2115 onwards.
@@
Another observation from Figure 2 is that SCC trajectories are very different across SSPs but almost the same between the 2050 and 2100 net-zero scenarios. Under SSP5, SCC is the largest achieving \$400 per ton in 2100, about \$150 under SSP3 and SSP4, about \$250 under SSP2 and about \$200 under SSP1. Note that SSP5 is the scenario of “Fossil fueled development” and thus leading to the largest social cost of carbon. In 2025, SCC is approximately \$30-50 across SSP markers and close to the original DICE results.
In addition to plots, the temperature results under the marker SSPs for the year 2100 are shown in Table 1. Here, we show results for the net-zero scenarios and zero industrial emission scenarios. There is almost no difference between net-zero and zero industrial emission scenarios. Repeating what we observed in Figure 2, the smallest temperature 2.5C is achieved under SSP1 and net-zero in 2050, and it is larger to 2.6C under other marker SSPs for net zero in 2050. For the 2100 net-zero scenario, SSPs lead to more variation in temperature ranging from 3C under SSP1 to 3.6C under SSP5.
Figure 2: DICE results for social cost of carbon ($SCC_{t}$, \$ per ton of $\mathrm{CO}_{2}$), temperature ($T_{t}^{\mathrm{AT}}$), and carbon concentration ($M_{t}^{\mathrm{AT}}$) under five marker SSPs under the scenarios achieving net-zero emissions in 2050 (red) and 2100 (blue). Black dashed line denotes the original DICE model results.
@@
For non-marker scenarios, we consider four IAMs: AIM/CGE, GCAM, IMAGE and WITCH-GLOBIOM that have results under all five SSPs in the SSPdata-IIASA database. Table 2 shows atmospheric temperature in the year 2100 produced by the DICE model calibrated to the non-marker scenarios. Here, we see results similar to Table 1. Net-zero and zero industry emission scenarios lead to approximately the same results. Under the 2050 net-zero scenario, temperature range is 2.5-2.7C across SSPs and IAMs. Under the 2100 net-zero scenario, temperature range is 3.0-3.7C. We also calculated trajectories of atmospheric temperature, atmospheric concentration and social cost of carbon under non-marker scenarios from 2015 to 2100 presented in Figures 3-6. These results are not materially different from the DICE results under the marker scenarios in Figure 2, confirming the consistency in predictions across considered process-based IAMs.

### v2

@@ 2.3 DICE calibration
Population $L_{t}$, total factor productivity $A_{t}$, land emission $E_{t}^{\mathrm{Land}}$, and decorbanisation $\sigma_{t}$ are exogenous parametric functions specified in DICE. In the original DICE, these were estimated using various data sources, resulting in a single scenario of the future for the economy and climate. In our study we estimate these functions from population, GDP and emission data of the baseline SSPs and process-based IAMs, to get many plausible scenarios to assess the possible trajectories of climate and economic variables in the future. Also note that, in our extended DICE model, $L_{t}$ is an endogenous state variable driven by global atmospheric temperature (2.1).
The SSPs provide quantitative projections and qualitative descriptions. Quantitative projections include population and economic growth. A detailed description of SSPs can be found in Riahi et al. (2017). In short, they are referred as follows.
  • SSP1 – “Sustainability - Taking the Green Road” (low challenges to mitigation and adaptation), representing the sustainable development.
  • SSP2 – “Middle of the Road” (medium challenges to mitigation and adaptation), implying a development pathway consistent with typical historical patterns.
  • SSP3 – “Regional Rivalry – A Rocky Road” (high challenges to mitigation and adaptation), characterised by international fragmentation and regional rivalry.
  • SSP4 – “Inequality – A Road Divided (low challenges to mitigation, high challenges to adaptation)”, characterised by extreme inequality both across and within countries.
  • SSP5 – “Fossil-fueled Development – Taking the Highway” (high challenges to mitigation, low challenges to adaptation), forecasting economic successes for both industrialized and emerging economies.
@@
Many process-based IAMs have been developed in climate change economics. These are detailed, multi-region, large-scale computer simulation models to assess policy options for stabilising the global climate. They couple representations of the economy, the energy system, the agricultural and land use system, and the climate system. The specific IAMs considered in our study will be described below. In principle, each process-based IAM can be run for each SSP. However, for reference, for each SSP a single IAM interpretation was selected as the so-called representative marker scenario for recommended use by future analyses of climate change, its impacts, and response measures. These are the so-called marker SSPs selected as representative of the broader developments of each SSP, guided by the internal consistency and the ability of the different models to represent distinct characteristics of the storylines. Below is a summary list of marker SSPs:
@@
For a detailed discussion, see also Riahi et al. (2017). Figure 1 shows the data for the world GDP, population, and emission under five marker SSPs (baselines). Here, we note that only under SSP3, the world population keeps growing from now to 2100. Under other SSPs, the population peaks at around 2050-2070. It is the lowest under SSP1 and the largest under SSP3. GDP grows over time under all SSPs; it is the lowest under SSP3 and the largest under SSP5. Industry emission grows under SSP2 and SSP3. It increases until 2050-2060 for SSP1 and SSP4, and then decreases. Under SSP5, emission is the largest, growing fast until around 2080 where it flattens. Under SSP1, emission is the lowest.
Figure 1: Global population (millions), GDP (trillions 2005 USD per year), and industrial $\mathrm{CO}_{2}$ emission (Gigatons) scenarios by marker SSPs (Red: SSP1, Grey: SSP2, Blue: SSP3, Green: SSP4, Brown: SSP5).
The SSP-IAM data are available from the SSPdata-IIASA database. In addition to data from marker SSPs, we also considered data for non-marker SSPs. Not all IAMs can run for all SSPs. For non-marker scenarios, we consider four IAMs: AIM/CGE, GCAM, IMAGE and WITCH-GLOBIOM Emmerling et al. (2016) that have results under all five SSPs.
We use the baseline SSP scenarios from the process-based IAMs to estimate the DICE exogenous functions as follows [^8] This approach is motivated by paper paper Yang et al. (2018) that examines the DICE results under the marker SSP1–SSP5, where the results of the China Climate Change integrated assessment model (C3IAM) were used to re-estimate some parameters in the DICE model. (the corresponding variables taken from the SSP-IAM scenario are indicated by $\widehat{}$ ).
  • The population $L_{t}$ is taken from the SSP-IAM scenario, i.e. $L_{t}=\widehat{L}_{t}$. Here we remark that, we tested to use $L_{t}$ as endogenous variable evolving in time according to (2.1), where death and fertility rates are estimated from SSP data but the impact of the excess mortality function $\delta(T^{\mathrm{AT}}_{t})$ on the temperature projections and other DICE outputs appeared to be not material, and thus one can simply use $L_{t}=\widehat{L}_{t}$.
  • The decarbonisation function $\sigma_{t}$ is estimated as $\sigma_{t}=\widehat{E}^{\mathrm{Ind}}_{t}$$\widehat{Y}_{t}$, where $\widehat{E}^{\mathrm{Ind}}_{t}$ is taken from the IAM baseline scenarios and $\widehat{Y}_{t}$ is the gross GDP in the baseline SSPs.
@@
  • The total productivity factor $A_{t}$ is calibrated so that the gross GDP $Y_{t}$ from DICE matches $\widehat{Y}_{t}$ the gross GDP in the baseline SSPs. This calibration involves iterative process that, for each iteration $m=1,2,$... calculates the DICE trajectories of $Y^{(m)}_{t}$ and model state variables $X^{(m)}_{t}$, t=0,1,,Nt=0,1,...,N using
@@
    until trajectory of $Y^{(m)}_{t}$ approximates the SSP trajectory of $\widehat{Y}_{t}$ closely. $K_{t}^{(m-1)}$, t=0,1,,Nt=0,1,...,N is calculated according to (3) and (4) via $A^{(m-1)}_{t}$, t=0,1,,Nt=0,1,...,N, i.e. iterative process (12) is the standard fixed-point iteration method. Here we note that: each iteration involves solving the DICE model (optimising the welfare function with respect to the consumption) without climate damage and abatement costs; $K^{(0)}_{t}$ is taken from the original DICE solution; and convergence is achieved quickly after few iterations.
Then for each baseline SSP and process-based IAM we solve DICE (miximising the welfare function with respect to consumption $C_{t}$) resulting in determinstic trajectory for temperature, carbon concentration, and social cost of carbon. We do not optimise welfare function in DICE for emission control $\widetilde{\mu}_{t}\in[0,1]$ but we consider scenarios when net-zero emission is achieved by a specific year as follows:
  • Net-zero in 2050: $\widetilde{\mu}_{t}$ increases linearly from 0 in 2015 to 1 in the year 2050 and remains 1 afterward; and
  • Net-zero in 2100: $\widetilde{\mu}_{t}$ increases linearly from 0 in 2015 to 1 in the year 2100 and remains 1 afterward.
@@ 3 Results
In this section, firstly we show results for climate-economy projections and then show impact of temperature-related excess mortality on two synthetic life portfolios of a life insurer.
3.1 DICE projections under SSPs and net-zero scenarios
Figure 2 shows the DICE results for the trajectories of atmospheric temperature $T_{t}^{\mathrm{AT}}$, atmospheric carbon concentration $M_{t}^{\mathrm{AT}}$, and social cost of carbon $SCC_{t}$ (USD per ton of $\mathrm{CO}_{2}$) under five SSP markers from 2015 to 2100. Each subplot in Figure 2 shows the trajectory under the original DICE (black dashed line), the trajectory under the scenario of achieving net-zero emissions in 2050 (red line), and the trajectory under the scenario of achieving net-zero emission in 2100 (blue line). Here, the original DICE trajectory is for reference. The original DICE trajectory corresponds to the optimal emission control and optimal consumption, and does not involve any SSP-IAM data calibration. A priori we do not know where the original DICE trajectory should be with respect to other scenarios plotted in Figure 2.
As expected we can see that the temperature under the scenario of net-zero in 2050 is lower than under the scenario of net-zero in 2100 across all SSPs. Under the 2050 net-zero scenario, the temperature in 2100 is about 2.5C for all SSPs while under the 2100 net-zero scenario, temperature in 2100 is just under 3C for SSP1 and materially larger than 3C for SSP2-SSP5 (highest under SSP5). This is of course not surprising because SSP1 corresponds to “Green road” scenario and SSP5 is “Fossil-fueled development” scenario. The original DICE trajectory of temperature is close 2100 net-zero scenario. This is somewhat expected because the optimal $\mu_{t}$ under original DICE reaches 1 in 2115 onwards.
@@ 3.1 DICE projections under SSPs and net-zero scenarios
Another observation from Figure 2 is that SCC trajectories are very different across SSPs but almost the same between the 2050 and 2100 net-zero scenarios. Under SSP5, SCC is the largest. In 2100, it is approximately USD 400 under SSP5, USD 150 under SSP3, USD 180 under SSP4, USD 250 under SSP2, and USD 200 under SSP1. Note that SSP5 is the scenario of “Fossil-fueled development” and thus it is not surprising to see it is leading to the largest social cost of carbon. It is interesting to note that the lowest SCC is under SSP3 where it is approximately USD 150 per ton in 2100 which is lower than under “Green road” SSP1 scenario. Careful calculation/consideration is required for $SCC_{t}$ as it represents the net present value of all economic damages along the future trajectory from the additional ton of $\mathrm{CO}_{2}$ released in year $t$. In 2025, SCC is approximately USD 30-50 across SSP markers and close to the original DICE results. These results for SCC under SSP markers are generally consistent with reported results in the literature, see e.g. Riahi et al. (2022) and Yang et al. (2018).
Figure 2: DICE results for social cost of carbon ($SCC_{t}$, USD per ton of $\mathrm{CO}_{2}$), temperature ($T_{t}^{\mathrm{AT}}$), and carbon concentration ($M_{t}^{\mathrm{AT}}$) under five marker SSPs under the scenarios achieving net-zero emissions in 2050 (red) and 2100 (blue). Black dashed line denotes the original DICE model results.
In addition to plots, the temperature results under the marker SSPs for the year 2100 are shown in Table 1. Here, we show results for the net-zero scenarios and zero industrial emission scenarios. There is almost no difference for temperature between net-zero and zero industrial emission scenarios. Repeating what we observed in Figure 2, the smallest temperature 2.5C is achieved under SSP1 and net-zero in 2050, and it is close to 2.6C under other marker SSPs for net zero in 2050. For the 2100 net-zero scenario, SSPs lead to more variation in temperature ranging from 3C under SSP1 to 3.7C under SSP5.
@@
For non-marker scenarios, we consider four IAMs: AIM/CGE, GCAM, IMAGE and WITCH-GLOBIOM that have results under all five SSPs in the SSPdata-IIASA database. Table 2 shows atmospheric temperature in the year 2100 produced by the DICE model calibrated to the non-marker scenarios. Here, we see results similar to Table 1. Net-zero and zero industry emission scenarios lead to approximately the same results. Under the 2050 net-zero scenario, temperature range is 2.5-2.7C across SSPs and IAMs. Under the 2100 net-zero scenario, temperature range is 3.0-3.7C. We also calculated trajectories of atmospheric temperature, atmospheric concentration and social cost of carbon under non-marker scenarios from 2015 to 2100 presented in Appendix A, Figures 5-8. These results are not materially different from the DICE results under the marker scenarios in Figure 2, confirming the consistency in predictions across considered process-based IAMs. Unfortunately, for all scenarios considered, we see that even under 2050 net-zero commitments, projected temperatures are in the range 2.5 – 2.7C by 2100 exceeding 1.5 – 2C target set out in the Paris Agreement, underscoring the need for negative emission technologies to be employed to achieve the target.
@@
3.2 Stress testing life insurance and annuity portfolios
In this subsection we use the mortality damage function derived in Bressler (2021) in conjunction with the Gompertz law of mortality as a baseline mortality law. We recall that the mortality damage function of Bressler (2021) is of the shape $\widetilde{\delta}(t)=at^{\nu}$ for $a=0.0001811$ and $\nu=3.745$. Moreover, the Gompertz law of mortality for a 25-year old takes the form $\lambda(t)=\frac{1}{b}\exp\left((25+t-M)/b\right)$, where the parameters $M$ = 88.23 and $b=9.38$ have been calibrated in Arandjelović et al. (2023) to mortality rates from many countries; for details, we refer to the appendix therein.
Figure 3: Projected percentage increase in climate-induced excess mortality rates under the DICE-calibrated marker SSPs and net zero scenarios. Each curve represents the mortality impact implied by the corresponding temperature trajectory between 2015 and 2100.
[…]
For numerical illustration of the impact of temperature-related excess mortality rate on life-contingent insurance products we consider a projected stress test scenario for two portfolios in the year 2100 with age distributions and amounts insured similar to Dong et al. (2024) and Tan and Bruhn (2019), see Figure 4:
  • Portfolio A: 10,000 life annuity products,
  • Portfolio B: 10,000 life insurance products.
For portfolio A, we assume that the insured amount is paid out at the end of the year, given that the client survives until then. We are interested in the distribution of the amount that is not paid out due to the fact that clients did not survive until the end of the year. Climate-induced excess mortality on this distribution is expected to increase the amount not paid due to the increase in mortality rate. Indeed, Table 3 confirms this rationale an demonstrates moderate but significant increase in the amounts not paid, both for the mean amount as well as in the tails of the distribution (1% and 99% percentiles). Similarly for portfolio B, where we are interested in the distribution of the total amount of insured sums paid out at the end of the year for every recorded client death, we can observe a noticeable increase in the mean amount and on the tails of the distribution.
Figure 4: Age distributions and corresponding total sum insured for the annuity and insurance portfolios. Bars show the number of clients at each age, while the dashed lines indicate the dollar value sum insured per client.
Table 3: Life annuity and life insurance portfolios under DICE-calibrated SSP scenarios.
Scenario Annuity Portfolio Insurance Portfolio
Mean $q_{0.01}$ $q_{0.99}$ Mean $q_{0.01}$ $q_{0.99}$
DICE 1.96 2.13 1.86 1.95 2.81 1.63
SSP1 2050 0.54 0.62 0.56 0.54 0.78 0.44
SSP1 2100 1.14 1.23 1.10 1.13 1.79 0.93
SSP2 2050 0.66 0.75 0.65 0.66 0.96 0.52
SSP2 2100 1.58 1.75 1.51 1.58 2.25 1.31
SSP3 2050 0.80 0.88 0.76 0.79 1.18 0.66
SSP3 2100 1.90 2.05 1.81 1.89 2.57 1.54
SSP4 2050 0.67 0.75 0.65 0.66 0.99 0.55
SSP4 2100 1.52 1.67 1.45 1.51 2.04 1.26
SSP5 2050 0.72 0.81 0.70 0.72 0.99 0.63
SSP5 2100 2.38 2.61 2.26 2.37 3.16 1.99
  • Values represent relative deviations (%) from the base case (no climate-induced impact on excess mortality) in mean, 1% percentile ($q_{0.01}$), and 99% percentile ($q_{0.99}$).
Finally, we check the impact of climate-induced excess mortality on human capital. We take the age-income function $y(s)$ from Arandjelović et al. (2023) and calculate the human capital of a 25-year old in the year 2015, defined as the net present value of future income,
$H=\int_{0}^{T}y(s)\mathrm{e}^{-rs-\int_{0}^{s}\widetilde{\lambda}(u)\,\mathrm{d}u}\,\mathrm{d}s,$ (13)
where $T=85$ denotes the maximal remaining lifetime, and $\widetilde{\lambda}(u)=\lambda(u)\times(1+f(u))$, with $f(u)$ being the scenario-dependent mortality damage function as in Figure 3, which has been approximated by a cubic polynomial for analytic tractability. Table 4 shows results. We can observe an almost negligible impact of climate-induced mortality on human capital, which can be explained as follows. As evident from Figure 3, the difference between projected mortality damages is almost negligible in the first half of this century, and only displays a more pronounced variability in the second half of the century. Due to discounting ($r=3.2\%$), the moderate impacts on mortality become almost negligible when expressed in terms of net present value.
Table 4: Human capital under DICE-calibrated marker SSP and net-zero scenarios. Values for $H$ are quoted in millions. $H_{\mathrm{rel}}$ denotes the relative deviation of $H$ from the base case in $\%$.
Scenario $H$ $H_{\mathrm{rel}}$
Base 1.4316 -
DICE 1.4314 -0.01
SSP1 2050 1.4315 -0.01
SSP1 2100 1.4314 -0.01
SSP2 2050 1.4315 -0.01
SSP2 2100 1.4314 -0.01
SSP3 2050 1.4314 -0.01
SSP3 2100 1.4314 -0.02
SSP4 2050 1.4315 -0.01
SSP4 2100 1.4314 -0.01
SSP5 2050 1.4315 -0.01
SSP5 2100 1.4313 -0.02